# Things we can learn by considering random locally symmetric manifolds

### v1

@@ 3 Invariant Random Subgroups
The results described in the previous section were established by considering the uniform measure on the finite set of $4$-regular graphs of a given size. This corresponds to a very primitive example of an IRS in $\text{SO}(n,1)$. We will now consider general IRS, which we will define for every locally compact second countable group. For a more comprehensive exposition of IRS we refer to the lecture note [Ge18a] (or [Ge18c]).
@@
The space $\operatorname{Sub}(G)$ is always compact but in general it is rather complicated. However it is a canonical $G$ space as $G$ acts on it by conjugation.
@@
Basic examples of IRS comes from normal subgroups and lattices:
@@
    A Dirac IRS, $\delta_{N}$ correspond to a closed normal subgroup $N\lhd G$.
@@
Remarkably, the study of IRS plays an important role in the development of the modern theory of lattices and their asymptotic invariants. We shall give several illustrations for this phenomena in subsequent sections.
IRS are strongly related to p.m.p actions. Indeed, given a probability space $(X,m)$ and a measure preserving $G$-action $G\curvearrowright X$, it follows by Varadarajan compact model that the stabilizer of almost every point in $X$ is closed in $G$. Thus the push-forward of $m$ under the stabilizer map $x\mapsto G_{x}$ is an IRS in $G$. By the following theorem every IRS arises in this way:
@@
This was proved in [AGV14] for discrete groups and in [7’s12, Theorem 2.4] for general $G$. The first thing that comes to mind is to take the given $G$ action on $(\text{Sub}(G),\mu)$, but then the stabilizer of a point $H\in\text{Sub}(G)$ is the normalizer $N_{G}(H)$ of $H$ rather than $H$. To correct this one considers the larger space $\text{Cos}_{G}$ of all cosets of all closed subgroups, as a measurable $G$-bundle over $\text{Sub}(G)$. Defining an appropriate invariant measure on $\text{Cos}_{G}\times\mathbb{R}$ and replacing each fiber by a Poisson process on it, gives the desired probability space.
@@
Let us now restrict to the case where $G$ is a semisimple Lie group. Treating IRS’s as a generalization of lattices, a natural thing to do is to establish IRS analogs for theorems about lattices. An important example is the Borel density theorem.
@@
By the Von-Neumann–Cartan theorem, every closed subgroup of $G$ is a Lie group. Consider the map $\psi:\operatorname{Sub}(G)\to\text{Gr}(\text{Lie}(G))$, $H\mapsto\text{Lie}(H)$. It is upper semi-continuous, hence measurable. Let $\nu=\psi_{*}(\mu)$. Then $\nu$ is $\text{Ad}(G)$-invariant and hence, by the Furstenberg lemma [F76], is supported on $\{\text{Lie}(G),\{0\}\}$. Finally observe that $\psi^{-1}(\text{Lie}(G))=G$ and $\psi^{-1}(\{0\})$ is the set of discrete subgroups of $G$. Assuming that $G$ is not an atom of $\mu$ the result about discreteness follows.
@@
Let us denote by $\operatorname{Sub}_{d}(G)$ the space of discrete subgroups of $G$. This is an open subset of the compact space $\operatorname{Sub}(G)$. We shall say that an IRS $\mu$ is discrete if $\mu(\operatorname{Sub}_{d}(G))=1$, and let $\text{IRS}_{d}(G)$ denote the space of discrete IRS of $G$. Since $G$ is isolated in $\operatorname{Sub}(G)$ (see [Z37], see [Ge18a]) we obtain the following useful fact:
@@ 4 Kazhdan Margulis theorem for IRS
There are by now three different proofs of Theorem 4.1. The original proof [KM68], a proof by Gromov [BGS85] which is more geometric and the proof from [Ge18b] which relies on IRS. Each of these proof has some advantages on the others. For example the original proof applies to all discrete subgroups (not just lattices) while the proof of Gromov applies to general Hadamard spaces (rather than symmetric spaces). The proof from [Ge18b] does not apply for general discrete groups but it does apply for general IRS for which it gives a stronger statement than the original theorem, even when restricting to lattices. However, the main advantage of that proof is in its simplicity. It happens occasionally in mathematics that a consideration of more general notions allows softer and less technical treatments. This depends however on formulating appropriate definitions.
@@
$\mu(\{\Gamma\in\text{Sub}_{G}:\Gamma\cap U_{\epsilon}\neq\{1\}\})<\epsilon$
@@
$m\{g\Gamma\in G/\Gamma:g\Gamma g^{-1}=\{1\}\}\geq 1/2$
@@
$K_{n}=\{\Gamma\in\text{Sub}_{G}:\Gamma\cap U_{n}=\{Id_{G}\}\}.$
@@ 5 The Stuck–Zimmer theorem
As follows from the Borel density theorem every transitive irreducible p.m.p action of a semisimple Lie group $G$ is of the form $G\curvearrowright G/\Gamma$ where $\Gamma\leq G$ is an irreducible lattice. Thus, in view of Theorem 3.2, Theorem 5.1 can be reformulated as follows:
@@
$G/\Gamma\to\operatorname{Sub}(G),\leavevmode\nobreak\ g\Gamma\mapsto gNg^{1},$
@@
Let $G$ be a locally compact group. A discrete subgroup $\Lambda\leq G$ is co-amenable if the Hilbert space $L_{2}(G/\Lambda)$ admits $G$ asymptotically invariant vectors.
@@ 6 An exotic IRS in rank one
In the lack of Margulis’ normal subgroup theorem there are IRS supported on non-lattices. Indeed, if $G$ has a lattice with an infinite index normal subgroup $N\lhd\Gamma$, arguing as in the previous section, one obtains an ergodic p.m.p. space for which almost any stabilizer is a conjugate of $N$.
@@ 7 The Benjamini–Schramm, the weak-* and the IRS topologies
Let $\mathcal{M}$ be the be the space of all (isometry classes of) pointed proper metric spaces equipped with the Gromov–Hausdorff topology. This is a huge space and for many applications it is enough to consider compact subspaces of it obtained by bounding the geometry. That is, let $f(\epsilon,r)$ be an integer valued function defined on $(0,1)\times{\mathbb{R}}^{>0}$, and let $\mathcal{M}_{f}$ consist of those spaces for which $\forall\epsilon,r$, the $\epsilon$-entropy of the $r$-ball $B_{X}(r,p)$ around the special point is bounded by $f(\epsilon,r)$, i.e. no $f(\epsilon,r)+1$ points in $B_{X}(r,p)$ form an $\epsilon$-discrete set. Then $\mathcal{M}_{f}$ is a compact subspace of $\mathcal{M}$.
In many situations one prefers to consider some variants of $\mathcal{M}$ which carry more information about the spaces. For instance when considering graphs, it may be useful to add colors and orientations to the edges. The Gromov–Hausdorff distance defined on these objects should take into account the coloring and orientation. Another example is smooth Riemannian manifolds, in which case it is better to consider framed manifolds, i.e. manifold with a chosen point and a chosen frame at the tangent space at that point. In that case, one replace the Gromov–Hausdorff topology by the ones determined by $(\epsilon,r)$ relations (see [7’s12, Section 3] for details), which remembers also the directions from the special point.
@@
It can be shown that the later map is one to one, and since $\text{IRS}_{d}(G)$ is compact, it is a homeomorphism to its image (see [7’s12, Corollary 3.4]). One can also characterize the image of this map which is the space of probability measures on $\mathcal{M}(X)$ which are invariant under the geodesic flow.
@@ 8 A result of the 7 Samurai
The idea behind Theorem 8.3 is that if $\mu_{n}$ are probability measures converging to a limit $\mu$ and $\mu$ is not ergodic, then there is an $L_{2}(\mu)$ function on $X$ which is $G$-invariant, orthogonal to the constants and with norm $1$. Approximating it one gets continuous function $f$ on $X$ which, as a function in $L_{2}(\mu)$ is almost $G$-invariant, orthogonal to the constants and with norm $1$. Thus for large $n$ we have that $f$ is almost invariant in $L_{2}(\mu_{n})$, almost orthogonal to the constants and with norm almost $1$. Since $G$ has property (T) it follows that there is an invariant $L_{2}(\mu_{n})$ function close to $f$, so $\mu_{n}$ cannot be ergodic.
@@
Thus, in order to prove Theorem 8.2 (equivalently 8.1) we have to exclude the cases $\mu=\delta_{G}$ and $\mu=\mu_{\Gamma}$. The case $\mu=\delta_{G}$ is impossible since $G$ is an isolated point in $\text{Sub}_{G}$ (see [Z37] or [Ge18a]). Let us now suppose that $\mu=\mu_{\Gamma}$ for some lattice $\Gamma\leq G$ and aim towards a contradiction. For this, we will adopt the formulation of Theorem 8.1 again. Thus we suppose that $M_{n}\to M=\Gamma\backslash X$ in the Benjamini–Schramm topology.
@@
This on itself does not contradict property (T) since the complement $M^{\prime\prime}=M\setminus M^{\prime}$ could (and actually has to) be very small (or even empty if $M$ is compact).
@@
Bearing in mind that $\text{vol}(M_{n})\to\infty$, we get that for large $n$, the complement $M_{n}^{\prime\prime}:=M_{n}\setminus M_{n}^{\prime}$ has arbitrarily large volume, and in particular $\text{vol}(M_{n}^{\prime})\leq\text{vol}(M_{n}^{\prime\prime})$. Now, this contradicts the assumption that $C$ is the Cheeger constant of $X$.∎
@@
As explained in the ‘alternative proof’ above, the argument using Cheeger constant could be replaced by local rigidity. However, property (T) is crucial in two other parts of the proof. Indeed, the proof relies on Theorem 8.3 as well as the Stuck–Zimmer theorem 5.1 (which in tern relies on property (T)).
@@
In [7’s12, 7’s17] Theorem 8.1 was proved in the greater generality of irreducible lattices in higher rank semisimple groups with property (T). In fact it is enough to suppose that $G$ is higher rank and at least one of the factors has property (T). In [Lev20] A. Levit proved the analog theorem for higher rank groups regardless of property (T) but assuming that the lattices $\Gamma_{n}$ are congruence (instead of Kazhdan’s property (T) he made use of Clozel’s property $\tau$). If Serre’s conjecture is true, then every higher rank irreducible lattice is congruence.
@@
Obviously the analog of Theorem 8.1 is false for $G=\text{SO}(n,1)$. For example, one can start with a compact hyperbolic manifold $M=\Gamma\backslash\mathcal{H}^{n}$ and take a descending sequence of finite index normal subgroups $N_{i}\lhd\Gamma$ such that the infinite intersection $\cap N_{i}$ is non-trivial and consider the covering manifolds $M_{i}=N_{i}\backslash\mathcal{H}^{n}$. Also the examples constructed in §2 have uniformly bounded injectivity radius. However, maybe for some ‘more rigid’ rank one groups (for instance, this is interesting for $\text{SP}(n,1)$) the analog of Theorem 8.1 holds for some mysterious reason.
@@ 9 Convergence of normalized Betti numbers
The most famous application of Theorem 8.1 is the convergence of normalized Betti numbers. This was first proved in [7’s11] for sequences of compact manifolds with injectivity radius bounded below (i.e. uniformly discrete sequences of uniform lattices) and in [ABBG23] for general sequences of manifolds. For the simplicity we shall restrict to the special case considered in [7’s11].
Suppose that $G$ is a simple Lie group with $\mathbb{R}$-rank at least $2$ and associated symmetric space $X=G/K$. We denote $\beta_{k}(X)$ the $k^{th}$ $L^{2}$-Betti number of $X$ (see also the analytic definition below). Then
@@
Now let $M=\Gamma\backslash X$ be a compact $X$-manifold. Let $\Delta_{k}$ be the Laplacian on differentiable $k$-forms on $M$. It is a symmetric, positive definite, elliptic operator with pure point spectrum. Write $e^{-t\Delta_{k}}(x,y)$ ($x,y\in M$) for the heat kernel on $k$-forms on $M$, then for each positive $t$ we have:
@@
Recall also that the usual Betti numbers of $M$ satisfies
@@
Finally, using that the usual Euler characteristic is equal to its $L^{2}$ analogue and that $\Delta_{k}^{(2)}$ has zero kernel if $k\neq\frac{1}{2}\dim X$ we derive Theorem 9.1.∎

### v2

@@ 3 Invariant Random Subgroups
The results described in the previous section were established by considering the uniform measure on the finite set of $4$-regular graphs of a given size. This corresponds to a very primitive example of an IRS in $\text{SO}(n,1)$. We will now consider general IRS, which we will define for every locally compact second countable group. For a more comprehensive exposition of IRS we refer to the lecture notes [Ge18a] (or [Ge18c]).
@@
The space $\operatorname{Sub}(G)$ is always compact but in general it is rather complicated. However, it is a canonical $G$ space as $G$ acts on it by conjugation.
@@
Basic examples of IRS come from normal subgroups and lattices:
@@
    A Dirac IRS, $\delta_{N}$ corresponds to a closed normal subgroup $N\lhd G$.
@@
Remarkably, the study of IRS plays an important role in the development of the modern theory of lattices and their asymptotic invariants. We shall give several illustrations for this phenomenon in subsequent sections.
IRS are strongly related to p.m.p actions. Indeed, given a probability space $(X,m)$ and a measure preserving $G$-action $G\curvearrowright X$, it follows by Varadarajan compact model [V63] that the stabilizer of almost every point in $X$ is closed in $G$. Thus the push-forward of $m$ under the stabilizer map $x\mapsto G_{x}$ is an IRS in $G$. By the following theorem every IRS arises in this way:
@@
This was proved in [AGV14] for discrete groups and in [7’s12, Theorem 2.4] for general $G$. The first thing that comes to mind is to take the given $G$ action on $(\text{Sub}(G),\mu)$, but then the stabilizer of a point $H\in\text{Sub}(G)$ is the normalizer $N_{G}(H)$ of $H$ rather than $H$. To correct this one considers the larger space $\text{Cos}_{G}$ of all cosets of all closed subgroups, as a measurable $G$-bundle over $\text{Sub}(G)$. Defining an appropriate invariant measure on $\text{Cos}_{G}\times\mathbb{R}$ and replacing each fiber by a Poisson process on it, gives the desired probability space (see [7’s12, §2] for details).
@@
Let us now restrict to the case where $G$ is a semisimple Lie group. Treating IRS as a generalization of lattices, a natural thing to do is to establish IRS analogs for theorems about lattices. An important example is the Borel density theorem.
@@
By the Von-Neumann–Cartan theorem, every closed subgroup of $G$ is a Lie group. Let $\text{Gr}(\text{Lie}(G))$ denote the Grassmanian variety of subspaces of the Lie algebra $\text{Lie}(G)$. Consider the map $\psi:\operatorname{Sub}(G)\to\text{Gr}(\text{Lie}(G))$, $H\mapsto\text{Lie}(H)$. It is upper semi-continuous, hence measurable. Let $\nu=\psi_{*}(\mu)$. Then $\nu$ is $\text{Ad}(G)$-invariant and hence, by the Furstenberg lemma [F76], is supported on $\{\text{Lie}(G),\{0\}\}$. Finally observe that $\psi^{-1}(\text{Lie}(G))=G$ and $\psi^{-1}(\{0\})$ is the set of discrete subgroups of $G$. Assuming that $G$ is not an atom of $\mu$ the result about discreteness follows.
@@
Let us denote by $\operatorname{Sub}_{d}(G)$ the space of discrete subgroups of $G$. By Cartan’s theorem, a closed subgroup of $G$ is a Lie group. Since the dimension is a semicontinuous function, the set $\operatorname{Sub}_{d}(G)$ of zero-dimensional subgroups is an open subset of the compact space $\operatorname{Sub}(G)$. We shall say that an IRS $\mu$ is discrete if $\mu(\operatorname{Sub}_{d}(G))=1$, and let $\text{IRS}_{d}(G)$ denote the space of discrete IRS of $G$. Since $G$ is isolated in $\operatorname{Sub}(G)$ (see [Z37], see [Ge18a]) we obtain the following useful fact:
@@ 4 Kazhdan Margulis theorem for IRS
There are by now three different proofs of Theorem 4.1. The original proof [KM68], a proof by Gromov [BGS85] which is more geometric and the proof from [Ge18b] which relies on IRS. Each of these proofs has some advantages on the others. For example, the original proof applies to all discrete subgroups (not just lattices) while the proof of Gromov applies to general Hadamard spaces (rather than symmetric spaces). The proof from [Ge18b] does not apply for general discrete groups but it does apply for general IRS for which it gives a stronger statement than the original theorem, even when restricting to lattices. However, the main advantage of that proof is its simplicity. It happens occasionally in mathematics that a consideration of more general notions allows softer and less technical treatments. This depends however on formulating appropriate definitions.
@@
$\mu(\{\Gamma\in\text{Sub}(G):\Gamma\cap U_{\epsilon}\neq\{1\}\})<\epsilon$
@@
$m(\{g\Gamma\in G/\Gamma:g\Gamma g^{-1}=\{1\}\})\geq 1/2$
@@
$K_{n}=\{\Gamma\in\text{Sub}_{G}:\Gamma\cap U_{n}=\{1\}\}.$
@@ 5 The Stuck–Zimmer theorem
As follows from the Borel density theorem [^5] Recall that the Borel density theorem says that if $G/H$ carries an invariant probability measure, then $H$ is Zariski dense. This imples that the connected component $H^{\circ}$ is normal in $G$ and hence a product of simple factors. When the action is non-trivial and irreducible, this implies that $H^{\circ}=\{1\}$. every non-trivial transitive irreducible p.m.p action of a semisimple Lie group $G$ is of the form $G\curvearrowright G/\Gamma$ where $\Gamma\leq G$ is an irreducible lattice. Thus, in view of Theorem 3.2, Theorem 5.1 can be reformulated as follows:
@@
$G/\Gamma\to\operatorname{Sub}(G),\leavevmode\nobreak\ g\Gamma\mapsto gNg^{-1},$
@@
Let $G$ be a locally compact group. A discrete subgroup $\Lambda\leq G$ is co-amenable if the Hilbert space $L^{2}(G/\Lambda)$ admits $G$ asymptotically invariant vectors.
@@ 6 An exotic IRS in rank one
In the lack of Margulis’ normal subgroup theorem, there are IRS supported on non-lattices. Indeed, if $G$ has a lattice with an infinite index normal subgroup $N\lhd\Gamma$, arguing as in the previous page, one obtains an ergodic p.m.p. space for which almost any stabilizer is a conjugate of $N$.
@@ 7 The Benjamini–Schramm, the weak-* and the IRS topologies
Let $\mathcal{M}$ be the space of all (isometry classes of) pointed proper metric spaces equipped with the Gromov–Hausdorff topology. This is a huge space and for many applications, it is enough to consider compact subspaces of it obtained by bounding the geometry. That is, let $f(\epsilon,r)$ be an integer-valued function defined on $(0,1)\times{\mathbb{R}}^{>0}$, and let $\mathcal{M}_{f}$ consist of those spaces for which $\forall\epsilon,r$, the $\epsilon$-entropy [^6] The maximal cardinality of an $\epsilon$-discrete set. of the $r$-ball $B_{X}(r,p)$ around the special point is bounded by $f(\epsilon,r)$, i.e. no $f(\epsilon,r)+1$ points in $B_{X}(r,p)$ form an $\epsilon$-discrete set. Then $\mathcal{M}_{f}$ is a compact subspace of $\mathcal{M}$.
In many situations, one prefers to consider some variants of $\mathcal{M}$ which carry more information about the spaces. For instance when considering graphs, it may be useful to add colors and orientations to the edges. The Gromov–Hausdorff distance defined on these objects should take into account the coloring and orientation. Another example is smooth Riemannian manifolds, in which case it is better to consider framed manifolds, i.e. manifold with a chosen point and a chosen frame at the tangent space at that point. In that case, one replaces the Gromov–Hausdorff topology with the ones determined by $(\epsilon,r)$ relations (see [7’s12, Section 3] for details), which remembers also the directions from the special point.
@@
It can be shown that the later map is one-to-one, and since $\text{IRS}_{d}(G)$ is compact, it is a homeomorphism to its image (see [7’s12, Corollary 3.4]). One can also characterize the image of this map which is the space of probability measures on $\mathcal{M}(X)$ which are invariant under the geodesic flow (see [AB22]).
@@ 8 A result of the 7 Samurai
The idea behind Theorem 8.3 is that if $\mu_{n}$ are probability measures converging to a limit $\mu$ and $\mu$ is not ergodic, then there is an $L^{2}(\mu)$ function on $X$ which is $G$-invariant, orthogonal to the constants and with norm $1$. Approximating it one gets continuous function $f$ on $X$ which, as a function in $L_{2}(\mu)$ is almost $G$-invariant, orthogonal to the constants, and with norm $1$. Thus for large $n$ we have that $f$ is almost invariant in $L_{2}(\mu_{n})$, almost orthogonal to the constants, and with norm almost $1$. Since $G$ has property (T) it follows that there is an invariant $L_{2}(\mu_{n})$ function close to $f$, so $\mu_{n}$ cannot be ergodic.
@@
Thus, in order to prove Theorem 8.2 (equivalently 8.1) we have to exclude the cases $\mu=\delta_{G}$ and $\mu=\mu_{\Gamma}$. The case $\mu=\delta_{G}$ is impossible since $G$ is an isolated point in $\text{Sub}(G)$ (see [Z37] or [Ge18a]). Let us now suppose that $\mu=\mu_{\Gamma}$ for some lattice $\Gamma\leq G$ and aim towards a contradiction. For this, we will adopt the formulation of Theorem 8.1 again. Thus we suppose that $M_{n}\to M=\Gamma\backslash X$ in the Benjamini–Schramm topology.
@@
This in itself does not contradict property (T) since the complement $M^{\prime\prime}=M\setminus M^{\prime}$ could (and actually has to) be very small (or even empty if $M$ is compact).
@@
Bearing in mind that $\text{vol}(M_{n})\to\infty$, we get that for large $n$, the complement $M_{n}^{\prime\prime}:=M_{n}\setminus M_{n}^{\prime}$ has arbitrarily large volume, and in particular $\text{vol}(M_{n}^{\prime})\leq\text{vol}(M_{n}^{\prime\prime})$. Now, this contradicts the assumption that $C$ is the Cheeger constant of $X$. See Figure 3. ∎
@@
As explained in the ‘alternative proof’ above, the argument using the Cheeger constant could be replaced by local rigidity. However, property (T) is crucial in two other parts of the proof. Indeed, the proof relies on Theorem 8.3 as well as the Stuck–Zimmer theorem 5.1 (which in turn relies on property (T)).
@@
In [7’s12, 7’s17] Theorem 8.1 was proved in the greater generality of irreducible lattices in higher rank semisimple groups with property (T). In fact it is enough to suppose that $G$ is higher rank and at least one of the factors has property (T). In [Lev20] A. Levit proved the analog theorem for higher rank groups regardless of property (T) but assuming that the lattices $\Gamma_{n}$ are congruence (instead of Kazhdan’s property (T) he made use of Clozel’s property $\tau$, see[Cl03]). If Serre’s conjecture is true, then every higher rank irreducible lattice is congruence.
@@
Obviously the analog of Theorem 8.1 is false for $G=\text{SO}(n,1)$. For example, one can start with a compact hyperbolic manifold $M=\Gamma\backslash\mathcal{H}^{n}$ and take a descending sequence of finite index normal subgroups $N_{i}\lhd\Gamma$ such that the infinite intersection $\cap N_{i}$ is non-trivial and consider the covering manifolds $M_{i}=N_{i}\backslash\mathcal{H}^{n}$. Also, the examples constructed in §2 have uniformly bounded injectivity radius. However, maybe for some ‘more rigid’ rank one groups (for instance, this is interesting for $\text{Sp}(n,1)$) the analog of Theorem 8.1 holds for some mysterious reason.
@@ 9 Convergence of normalized Betti numbers
The most famous application of Theorem 8.1 is the convergence of normalized Betti numbers. This was first proved in [7’s11] for sequences of compact manifolds with injectivity radius bounded below (i.e. uniformly discrete sequences of uniform lattices) and in [ABBG23] for general sequences of manifolds. For simplicity we shall restrict to the special case considered in [7’s11].
Suppose that $G$ is a simple Lie group with $\mathbb{R}$-rank at least $2$ and associated symmetric space $X=G/K$. Let $X^{*}$ denote the compact dual of $X$ and recall that the Killing form on $\text{Lie}(G)$ induces a Riammanian structure and a volume form on $X^{*}$ as well as on $X$. We denote $\beta_{k}(X)$ the $k^{th}$ $L^{2}$-Betti number of $X$ (see also the analytic definition below). Then
@@
Now let $M=\Gamma\backslash X$ be a compact $X$-manifold. Let $\Delta_{k}$ be the Laplacian on differentiable $k$-forms on $M$. It is a symmetric, positive definite, elliptic operator with a pure point spectrum. Write $e^{-t\Delta_{k}}(x,y)$ ($x,y\in M$) for the heat kernel on $k$-forms on $M$, then for each positive $t$ we have:
@@
By Hodge’s theory, the usual Betti numbers of $M$ satisfies
@@
(Intuitively, the solutions of the heat equations converge to solutions of the Laplace equations when the time goes to infinity — the temperature stabilizes with time.)
@@
Finally, using that the usual Euler characteristic is equal to its $L^{2}$ analog and that $\Delta_{k}^{(2)}$ has zero kernel if $k\neq\frac{1}{2}\dim X$ we derive Theorem 9.1.∎